# Eugene Kritchevski

### v1

@@ Abstract
We introduce an elementary framework that captures the mechanism driving the alignment of rows and columns in products of positive matrices. All worst-case misalignment occurs already in dimension two, leading to an explicit collapse principle and a sharp nonlinear bound for finite products. The proof avoids Hilbert-metric and cone-theoretic techniques, relying instead on basic calculus. In the Hilbert metric, the classical Birkhoff–Bushell contraction captures only the linearized asymptotic regime, whereas our nonlinear envelope function gives the exact worst-case behavior for finite products.
@@ 1 Introduction
Products of positive matrices exhibit a pronounced alignment phenomenon: as the product grows, all rows and columns become nearly proportional. Classical methods approach this behavior through the Hilbert projective metric, probabilistic tools, or spectral theory. In this work we take a different route: we work directly with the distortion
$R(A):=\max_{i,j,k,\ell}\frac{a_{ik}a_{j\ell}}{a_{i\ell}a_{jk}},\qquad A\in\mathbb{R}_{>0}^{d_{1}\times d_{2}},$
which quantifies how far the rows or columns of $A$ are from being proportional. We have $R(A)=1$ in the perfectly aligned case; in general $R(A)\geq 1$ and increases with misalignment. Although $R(A)$ corresponds to the exponential of the projective diameter of $A$, we rely only on its elementary ratio form and use no projective-metric theory.
Tracking how distortion changes under matrix multiplication gives an elementary, coordinate-level mechanism for alignment. The central observation is that worst–case distortion occurs already in dimension two; higher dimensions do not produce any new extremal behaviour. In dimension two, the explicit sharp distortion bound can then be computed by an elementary one–variable calculus optimization. A dimension collapse argument shows that exactly the same bound holds in arbitrary dimension. This yields the envelope inequality
$R(AB)\leq\Phi(R(A),R(B)),$
where
$\Phi(\alpha,\beta)=\left(\frac{1+\sqrt{\alpha\beta}}{\sqrt{\alpha}+\sqrt{\beta}}\right)^{\!2}.$
Roadmap.
Section 2 introduces distortion and proves the envelope inequality in dimension 2. Section 3 extends the envelope inequality to arbitrary dimension via a collapse argument. Section 4 develops distortion for longer products of matrices, including explicit Frobenius–Perron bounds. Section 5 compares our result with the classical Birkhoff–Bushell theory. Section 6 highlights several directions for further work. Additional calculus lemmas used in the proofs are collected in the Appendix.
2 Distortion Calculus in Dimension 22
@@
Both $F(A)$ and $R(A)$ measure how far the two columns (equivalently, the two rows) of $A$ are from being proportional. Perfect proportionality is equivalent to $F(A)=R(A)=1$. The following properties are immediate from the definition:
  • Scaling invariance. Multiplying any row or any column of $A$ by a positive scalar leaves both $F(A)$ and $R(A)$ unchanged.
  • Effect of swapping. Swapping the two rows or the two columns of $A$ replaces $F(A)$ by $1/F(A)$ and therefore leaves $R(A)$ unchanged.
@@
    In the perfectly aligned case, $\det(A)=0$ and both orientations coincide. Note that in our projective, scale-invariant setting, the area interpretation of the determinant is irrelevant; only the sign information matters.
Our goal is to obtain a sharp explicit upper bound on the distortion
$R(AB)$
of the product of two positive $2\times 2$ matrices. The analysis is greatly simplified by a few distortion-preserving normalizations.
We first ensure positive orientation:
@@
so that $F(A)=R(A)$ and $F(B)=R(B)$. This can always be achieved by swapping the rows of $A$ and/or the columns of $B$; such swaps do not affect the distortion of $A$, $B$, or $AB$. After ensuring positive orientation of $A$ and $B$, we have $\det(A)\geq 0$, $\det(B)\geq 0$, and so $\det(AB)=\det(A)\det(B)\geq 0$. Then $F(AB)\geq 1$ and $AB$ is also positively oriented: $F(AB)=R(AB)$.
Next, we are free to multiply the rows of $A$ and the columns of $B$ by any positive scalars. Such scalings affect neither the orientation nor the distortion of $A$, $B$, or $AB$.
After these normalizations, we may assume, without loss of generality, that $A$ and $B$ are in the canonical form
@@
A direct computation gives
$AB=\begin{pmatrix}1+uv&1+\beta\,uv\\[5.69054pt] 1+\alpha\,uv&1+\alpha\beta\,uv\end{pmatrix}=\begin{pmatrix}1+t&1+\beta t\\[5.69054pt] 1+\alpha t&1+\alpha\beta t\end{pmatrix}$,
@@
$t=uv>0.$
Then
$R(AB)=f_{\alpha,\beta}(uv),$
where
$f_{\alpha,\beta}(t):=\frac{(1+t)(1+\alpha\beta t)}{(1+\beta t)(1+\alpha t)}\qquad t>0.$
Thus the problem of finding a sharp upper bound for $R(AB)$ in terms of $R(A)$ and $R(B)$ reduces to a simple calculus exercise: it suffices to maximize the function $f_{\alpha,\beta}(t)$ for $t>0$. The details are carried out in Lemma A.1 and the result is
$\max_{t>0}f_{\alpha,\beta}(t)=\Phi(\alpha,\beta),$
where
$\Phi(\alpha,\beta)=\frac{\bigl(1+\sqrt{\alpha\beta}\bigr)^{2}}{\bigl(\sqrt{\alpha}+\sqrt{\beta}\bigr)^{2}}.$
We refer to $\Phi$ as the envelope function.
Thus we obtain
$R(AB)=f_{\alpha,\beta}(uv)\leq\Phi(\alpha,\beta),$
which proves the following bound:
@@
$\boxed{R(AB)\leq\Phi\bigl(R(A),R(B)\bigr)=\left(\frac{1+\sqrt{R(A)R(B)}}{\sqrt{R(A)}+\sqrt{R(B)}}\right)^{\!2}}.$
Moreover, the envelope function $\Phi(\alpha,\beta)$ is increasing in each variable (Lemma A.2). Therefore, from Theorem 2.2 we obtain:
@@
Let $A$ and $B$ be strictly positive $2\times 2$ matrices, and suppose
$R(A)\leq\alpha,\qquad R(B)\leq\beta.$
Then
@@
We note that the envelope inequality is not merely an upper bound; the maximal distortion is attainable.
@@
    Given any $A>0$ with $R(A)=\alpha$, one can choose $B>0$ with $R(B)=\beta$ such that
    $R(AB)=\Phi(\alpha,\beta).$
@@
    Given any $B>0$ with $R(B)=\beta$, one can choose $A>0$ with $R(A)=\alpha$ such that
    $R(AB)=\Phi(\alpha,\beta).$
@@
We first prove (i). By Lemma A.1, the function $f_{\alpha,\beta}(t)$ attains its maximum $\Phi(\alpha,\beta)$ at
$t^{*}=\frac{1}{\sqrt{\alpha\beta}}.$
Every product of the form
$\begin{pmatrix}1&u\\[5.69054pt] 1&\alpha u\end{pmatrix}\begin{pmatrix}1&1\\[5.69054pt] v&\beta v\end{pmatrix},\qquad uv=t^{*},$
@@
Let $A>0$ be given with $R(A)=\alpha$. By rescaling the rows of $A$ by positive factors (which does not change $R(A)$ nor $R(AB)$), we may assume that
$A=\begin{pmatrix}1&x\\[5.69054pt] 1&y\end{pmatrix},\qquad x,y>0.$
If $A$ is positively oriented, i.e. $y/x=\alpha$, we take
$B=\begin{pmatrix}1&1\\[5.69054pt] v&\beta v\end{pmatrix},\qquad v=\frac{t^{*}}{x}.$
If $A$ is negatively oriented, i.e. $x/y=\alpha$, we take
$B=\begin{pmatrix}1&1\\[5.69054pt] \beta v&v\end{pmatrix},\qquad v=\frac{t^{*}}{y}.$
In both cases one checks directly that
$R(AB)=\Phi(\alpha,\beta).$
This proves (i).
For (ii), use that $R(M)=R(M^{\mathsf{T}})$ and $\Phi(\alpha,\beta)=\Phi(\beta,\alpha)$. Given $B$ with $R(B)=\beta$, apply (i) to $B^{\mathsf{T}}$ and $\alpha$: there exists $C>0$ with $R(C)=\alpha$ such that
@@
$\boxed{\Phi(\alpha,\beta)=\max\bigl\{\,R(AB):A,B>0,\ R(A)\leq\alpha,\ R(B)\leq\beta\,\bigr\}}.$
3 Reduction to Two Dimensions in the General Case
The goal of this section is to extend the envelope inequality to arbitrary dimension.
@@ 3 Reduction to Two Dimensions in the General Case
$s^{+}(x,y)=\max_{i}\frac{y_{i}}{x_{i}},\qquad s^{-}(x,y)=\min_{i}\frac{y_{i}}{x_{i}},$
and the distortion
$\operatorname{Dist}(x,y)=\frac{s^{+}(x,y)}{s^{-}(x,y)}.$
For perfectly aligned vectors we have $s^{+}(x,y)=s^{-}(x,y)$, so $\operatorname{Dist}(x,y)=1$. When the ratios $y_{i}/x_{i}$ are more scattered, the distortion increases. Figure 1 illustrates $s^{+}(x,y)$ and $s^{-}(x,y)$ as maximal and minimal slopes of the points $(x_{i},y_{i})$ in $\mathbb{R}_{>0}^{2}$.
Figure 1: The pair $(x,y)$ represented as a cloud of points between the extremal rays of slopes $s^{+}(x,y)$ and $s^{-}(x,y)$ in $\mathbb{R}_{>0}^{2}$.
@@
$R(A)=\max_{\begin{subarray}{c}1\leq i,j\leq d_{1}\\ 1\leq k,\ell\leq d_{2}\end{subarray}}\frac{a_{ik}\,a_{j\ell}}{a_{i\ell}\,a_{jk}}.$
Thus $R(A)$ is the maximum of the distortions of all $2\times 2$ submatrices of $A$, obtained by choosing any two rows and any two columns. Equivalently, it is the maximum distortion between any two columns (or two rows) of $A$:
@@
$F_{d}(x,y;u,v):=\frac{(x\cdot u)(y\cdot v)}{(x\cdot v)(y\cdot u)}.$
Here $x\cdot u$ denotes the standard Euclidean dot product.
@@
We will first construct $y^{*}$ and then use a similar procedure to obtain $v^{*}$. The strategy is to adjust the coordinates of $y$ one at a time, always modifying them in a way that does not decrease the value of $F_{d}(x,y;u,v)$, and while keeping the extremal slopes $s^{\pm}(x,y)$ unchanged.
To make this precise, partition the index set $\{1,\dots,d\}$ into
$I_{-}:=\{\,i:y_{i}/x_{i}=a^{-}\,\},\qquad I_{+}:=\{\,i:y_{i}/x_{i}=a^{+}\,\},$
@@
$I_{0}:=\{1,\dots,d\}\setminus(I_{-}\cup I_{+}).$
We will never alter the coordinates in $I_{-}$ or $I_{+}$, so that the extremal slopes remain fixed at all times. Thus we immediately set
$y_{i}^{*}:=y_{i}\qquad\text{for all }i\in I_{-}\cup I_{+}.$
The coordinatewise replacement procedure is applied only to indices in $I_{0}$.
We begin with a simple coordinate monotonicity observation. Fix $x,u,v$ and all coordinates of $y$ except $y_{i}$. Then
$f(y_{i}):=F_{d}(x,y;u,v)$
is a fractional linear function of $y_{i}$ with strictly positive denominator, and is therefore monotone (either increasing or decreasing) in $y_{i}$. If $f$ is increasing in $y_{i}$, we set
$y_{i}^{*}:=a^{+}\,x_{i},$
while if $f$ is decreasing in $y_{i}$, we set
$y_{i}^{*}:=a^{-}\,x_{i}.$
@@
$F_{d}(x,y^{*};u,v)\geq F_{d}(x,y;u,v),\qquad s^{\pm}(x,y^{*})=s^{\pm}(x,y),$

### v2

@@ Abstract
We study the deviation from proportionality of rows and columns in products of positive matrices. We prove a sharp, dimension-free bound showing that worst-case misalignment is already captured in dimension two and follows an explicit Möbius law, refining the classical Birkhoff–Bushell contraction theory.
@@ 1 Introduction
Products of positive matrices exhibit a pronounced alignment phenomenon: as the product grows, all columns (respectively, all rows) become nearly proportional. While alignment is well understood asymptotically, the finite-time law governing worst-case misalignment has not yet been isolated.
In this paper, we develop an elementary framework based on the distortion functional
$R(A):=\max_{\begin{subarray}{c}1\leq i,j\leq d_{1}\\ 1\leq k,\ell\leq d_{2}\end{subarray}}\frac{a_{ik}\,a_{j\ell}}{a_{i\ell}\,a_{jk}},\qquad A\in\mathbb{R}_{>0}^{d_{1}\times d_{2}}.$
Our main result is a sharp envelope inequality controlling the distortion of matrix products,
$R(AB)\leq\Phi(R(A),R(B)),\qquad\Phi(\alpha,\beta):=\left(\frac{1+\sqrt{\alpha\beta}}{\sqrt{\alpha}+\sqrt{\beta}}\right)^{2}.$
A key structural observation is a rigidity phenomenon for projective distortion: extremal misalignment does not engage genuinely high-dimensional geometry, but collapses to a two-dimensional extremal configuration.
While distortion is classically measured by the projective diameter $\log R(A)$ in the Hilbert metric, we show that $\sqrt{R(A)}$ is the natural scale on which extremal distortion evolves according to an explicit Möbius law.
Organization of the paper.
Section 2 establishes the envelope inequality in dimension 2. Section 3 extends the result to arbitrary dimension via a dimension-reduction argument. Section 4 derives explicit finite-time bounds for the distortion of matrix products. Section 5 compares the envelope bound with the classical Birkhoff–Bushell theory.
2 Distortion Bounds in Dimension 22
@@
Both $F(A)$ and $R(A)$ measure the deviation of the two columns (or rows) of $A$ from proportionality. Perfect proportionality is equivalent to $F(A)=R(A)=1$. The following properties are immediate.
  • Scaling invariance. Multiplying any row or column of $A$ by a positive scalar leaves $F(A)$ and $R(A)$ unchanged.
  • Effect of swapping. Swapping the two rows or the two columns replaces $F(A)$ by $1/F(A)$ and leaves $R(A)$ unchanged.
@@
    In the proportional case $\det(A)=0$, both orientations coincide.
Our goal is to obtain a sharp explicit upper bound on $R(AB)$ for the product of two positive $2\times 2$ matrices. The analysis is simplified by distortion-preserving normalizations.
We first ensure positive orientation,
@@
so that $F(A)=R(A)$ and $F(B)=R(B)$. This can be achieved by swapping rows of $A$ and/or columns of $B$, which does not affect the distortions of $A$, $B$, or $AB$. Then $\det(A),\det(B)\geq 0$, and hence $\det(AB)\geq 0$, so that $F(AB)=R(AB)$.
We may also multiply the rows of $A$ and the columns of $B$ by arbitrary positive scalars. Such scalings affect neither the orientation nor the distortion of $A$, $B$, or $AB$.
After these normalizations, we may assume without loss of generality that
@@
Then
$AB=\begin{pmatrix}1+t&1+\beta t\\[5.69054pt] 1+\alpha t&1+\alpha\beta t\end{pmatrix},\qquad t=uv>0.$
Accordingly,
$R(AB)=f_{\alpha,\beta}(t),\qquad f_{\alpha,\beta}(t):=\frac{(1+t)(1+\alpha\beta t)}{(1+\beta t)(1+\alpha t)},\quad t>0.$
The task of bounding $R(AB)$ therefore reduces to maximizing the one-variable function $f_{\alpha,\beta}(t)$. Differentiating its logarithm yields
$\frac{d}{dt}\log f_{\alpha,\beta}(t)=\frac{(\alpha-1)(\beta-1)\,(1-\alpha\beta t^{2})}{(1+t)(1+\alpha t)(1+\beta t)(1+\alpha\beta t)}.$
Thus $f_{\alpha,\beta}$ is maximized at
$t^{*}:=\frac{1}{\sqrt{\alpha\beta}},$
with
$f_{\alpha,\beta}(t^{*})=\Phi(\alpha,\beta),$
@@
$\Phi(\alpha,\beta):=\left(\frac{1+\sqrt{\alpha\beta}}{\sqrt{\alpha}+\sqrt{\beta}}\right)^{2}.$
We refer to $\Phi(\alpha,\beta)$ as the envelope function. Hence
$R(AB)\leq\Phi(\alpha,\beta),$
and we obtain the following result.
@@
$R(AB)\leq\Phi\bigl(R(A),R(B)\bigr)=\left(\frac{1+\sqrt{R(A)R(B)}}{\sqrt{R(A)}+\sqrt{R(B)}}\right)^{\!2}.$
The envelope function $\Phi$ is increasing in each variable, yielding the following corollary.
@@
Let $A$ and $B$ be strictly positive $2\times 2$ matrices with $R(A)\leq\alpha$ and $R(B)\leq\beta$. Then
@@
The envelope inequality is sharp.
@@
    For any $A>0$ with $R(A)=\alpha$, there exists $B>0$ with $R(B)=\beta$ such that $R(AB)=\Phi(\alpha,\beta)$.
@@
    For any $B>0$ with $R(B)=\beta$, there exists $A>0$ with $R(A)=\alpha$ such that $R(AB)=\Phi(\alpha,\beta)$.
@@
We first prove (i). Every product of the form
$\begin{pmatrix}1&u\\ 1&\alpha u\end{pmatrix}\begin{pmatrix}1&1\\ v&\beta v\end{pmatrix},\qquad uv=t^{*},$
@@
Let $A>0$ satisfy $R(A)=\alpha$. By rescaling rows, we may write
$A=\begin{pmatrix}1&x\\ 1&y\end{pmatrix},\qquad x,y>0.$
If $A$ is positively oriented, so that $y/x=\alpha$, we take
$B=\begin{pmatrix}1&1\\ v&\beta v\end{pmatrix},\qquad v=\frac{t^{*}}{x}.$
If $A$ is negatively oriented, so that $x/y=\alpha$, we take
$B=\begin{pmatrix}1&1\\ \beta v&v\end{pmatrix},\qquad v=\frac{t^{*}}{y}.$
In both cases, $R(AB)=\Phi(\alpha,\beta)$, proving (i).
For (ii), we use the symmetry under transposition. Note that
$R(M)=R(M^{\mathsf{T}})\qquad\text{and}\qquad\Phi(\alpha,\beta)=\Phi(\beta,\alpha).$
Given $B>0$ with $R(B)=\beta$, apply part (i) to $B^{\mathsf{T}}$. There exists $C>0$ with $R(C)=\alpha$ such that
@@
$\Phi(\alpha,\beta)=\max\bigl\{R(AB):A,B>0,\ R(A)\leq\alpha,\ R(B)\leq\beta\bigr\}.$
3 Reduction to Two Dimensions
In this section we extend the notion of distortion to arbitrary dimension and prove the envelope inequality via a reduction to two dimensions.
@@ 3 Reduction to Two Dimensions
$s^{+}(x,y):=\max_{i}\frac{y_{i}}{x_{i}},\qquad s^{-}(x,y):=\min_{i}\frac{y_{i}}{x_{i}}.$
Figure 1 illustrates $s^{\pm}(x,y)$ as the extremal slopes of the points $(x_{i},y_{i})$ in $\mathbb{R}_{>0}^{2}$.
Figure 1: Geometric representation of $(x,y)$ as points in $\mathbb{R}_{>0}^{2}$.
The misalignment of $x$ and $y$ is measured by the distortion
$\operatorname{Dist}(x,y):=\frac{s^{+}(x,y)}{s^{-}(x,y)}.$
Equivalently,
$\operatorname{Dist}(x,y)=\max_{i,j}\frac{y_{i}x_{j}}{y_{j}x_{i}}.$
@@
$R(A):=\max_{\begin{subarray}{c}1\leq i,j\leq d_{1}\\ 1\leq k,\ell\leq d_{2}\end{subarray}}\frac{a_{ik}\,a_{j\ell}}{a_{i\ell}\,a_{jk}}.$
Thus $R(A)$ is the maximum distortion of all $2\times 2$ submatrices of $A$. Equivalently,
@@
The following functional encodes the oriented distortion of the $2\times 2$ block of a matrix product, obtained by pairing rows $x,y$ of the left factor with columns $u,v$ of the right factor.
@@
$F_{d}(x,y;u,v):=\frac{(x\cdot u)(y\cdot v)}{(x\cdot v)(y\cdot u)},$
where $\cdot$ denotes the standard Euclidean dot product.
@@
We first construct $y^{*}$. The strategy is to use monotonicity to adjust the coordinates of $y$ one at a time without decreasing the value of $F_{d}$.
Define
$f(z_{1},\dots,z_{d})\;:=\;F_{d}(x,z;u,v),\qquad z=(z_{1},\dots,z_{d})\in\mathbb{R}_{>0}^{d}.$
For $i\in\{1,\dots,d\}$, write $z_{\neq i}:=(z_{1},\dots,z_{i-1},z_{i+1},\dots,z_{d})$ and
$f_{i}(t\mid z_{\neq i})\;:=\;f(z_{1},\dots,z_{i-1},t,z_{i+1},\dots,z_{d}),\qquad t>0.$
For fixed $z_{\neq i}$, the function $t\mapsto f_{i}(t\mid z_{\neq i})$ is fractional linear with strictly positive denominator, and hence is either increasing, decreasing, or constant.
Set
$B:=\{\,z\in\mathbb{R}_{>0}^{d}:a^{-}x_{i}\leq z_{i}\leq a^{+}x_{i}\text{ for all }i\,\}.$
For each $i\in\{1,\dots,d\}$, define the coordinate update map $E_{i}$ on $B$ by
[…]
Then
$E_{i}(z)\in\{a^{-}x_{i},a^{+}x_{i}\}\quad\text{for all }z\in B.$
Define the associated vector update operator $\mathcal{E}_{i}:B\to B$ by
$(\mathcal{E}_{i}z)_{j}:=\begin{cases}z_{j},&j\neq i,\\ E_{i}(z),&j=i.\end{cases}$
By construction,
[…]
Define
$y^{*}\;:=\;(\mathcal{E}_{d}\circ\cdots\circ\mathcal{E}_{1})(y).$
Then
$f(y^{*})\;\geq\;f(y),\qquad s^{\pm}(x,y^{*})\;=\;s^{\pm}(x,y),\qquad\frac{y_{i}^{*}}{x_{i}}\in\{a^{-},a^{+}\}\ \ \text{for all }i.$
We next construct $v^{*}$ in the same manner. Define
$g(w_{1},\dots,w_{d})\;:=\;F_{d}(x,y^{*};u,w),\qquad w=(w_{1},\dots,w_{d})\in\mathbb{R}_{>0}^{d},$
@@
$g_{i}(t\mid w_{\neq i})\;:=\;g(w_{1},\dots,w_{i-1},t,w_{i+1},\dots,w_{d}),\qquad t>0.$
For fixed $w_{\neq i}$, the function $t\mapsto g_{i}(t\mid w_{\neq i})$ is fractional linear with strictly positive denominator. Repeating the coordinate update argument in this setting produces $v^{*}\in\mathbb{R}_{>0}^{d}$ such that
$g(v^{*})\;\geq\;g(v),\qquad s^{\pm}(u,v^{*})\;=\;s^{\pm}(u,v),\qquad\frac{v_{i}^{*}}{u_{i}}\in\{b^{-},b^{+}\}\ \ \text{for all }i.$
Finally, by construction
$F_{d}(x,y^{*};u,v^{*})$ $=g(v^{*})$
$\geq g(v)$
$=F_{d}(x,y^{*};u,v)$
$=f(y^{*})$
$\geq f(y)$
$=F_{d}(x,y;u,v),$
which completes the proof. ∎
The following lemma shows that extremal configurations collapse to dimension two.
Lemma 3.2 (Two-dimensional collapse).
Let $x,y,u,v\in\mathbb{R}_{>0}^{d}$ and set
$a^{\pm}:=s^{\pm}(x,y),\qquad b^{\pm}:=s^{\pm}(u,v).$
Assume that
$\frac{y_{i}}{x_{i}}\in\{a^{-},a^{+}\}\quad\text{and}\quad\frac{v_{i}}{u_{i}}\in\{b^{-},b^{+}\}\qquad\text{for all }1\leq i\leq d.$
Then there exist vectors $x^{\prime},y^{\prime},u^{\prime},v^{\prime}\in\mathbb{R}_{>0}^{2}$ such that
$s^{\pm}(x^{\prime},y^{\prime})=a^{\pm},\qquad s^{\pm}(u^{\prime},v^{\prime})=b^{\pm},$
and
$F_{2}(x^{\prime},y^{\prime};u^{\prime},v^{\prime})\;\geq\;F_{d}(x,y;u,v).$
Proof.
If $a^{-}=a^{+}=a$, then $y=ax$ and $F_{d}(x,y;u,v)=1$. In this case one may take $x^{\prime}=(1,1)$, $y^{\prime}=(a,a)$, $u^{\prime}=(1,1)$ and $v^{\prime}=(b^{-},b^{+})$, so that $s^{\pm}(x^{\prime},y^{\prime})=a$, $s^{\pm}(u^{\prime},v^{\prime})=b^{\pm}$, and $F_{2}(x^{\prime},y^{\prime};u^{\prime},v^{\prime})=1$. The case $b^{-}=b^{+}$ is similar and is handled analogously. In either case the conclusion of the lemma holds.
We therefore assume from now on that $a^{-}<a^{+}$ and $b^{-}<b^{+}$.
Define the index sets
$I_{-}:=\{\,i:\frac{y_{i}}{x_{i}}=a^{-}\,\},\qquad I_{+}:=\{\,i:\frac{y_{i}}{x_{i}}=a^{+}\,\},$
which form a partition of $\{1,\dots,d\}$ and are both nonempty by assumption.
Step 1: Two–dimensional reduction preserving $F_{d}$.
Define
$x^{-}:=\sum_{i\in I_{-}}x_{i},\qquad x^{+}:=\sum_{i\in I_{+}}x_{i},\qquad x^{\prime}:=(x^{-},x^{+})\in\mathbb{R}_{>0}^{2},$
and
$u^{-}:=\frac{1}{x^{-}}\sum_{i\in I_{-}}x_{i}u_{i},\qquad u^{+}:=\frac{1}{x^{+}}\sum_{i\in I_{+}}x_{i}u_{i},\qquad u^{\prime}:=(u^{-},u^{+})\in\mathbb{R}_{>0}^{2}.$
Set
$y^{\prime}:=(a^{-}x^{-},\,a^{+}x^{+})\in\mathbb{R}_{>0}^{2}.$
By construction, $s^{\pm}(x^{\prime},y^{\prime})=a^{\pm}$. A direct computation shows that
$x\cdot u=x^{\prime}\cdot u^{\prime},\qquad y\cdot u=y^{\prime}\cdot u^{\prime}.$
Define an auxiliary vector $\tilde{v}\in\mathbb{R}_{>0}^{2}$ by
[…]
@@
$x\cdot v=x^{\prime}\cdot\tilde{v},\qquad y\cdot v=y^{\prime}\cdot\tilde{v}$,
and hence
[…]
Let
$\tilde{b}^{-}:=\frac{\tilde{v}^{-}}{u^{-}},\qquad\tilde{b}^{+}:=\frac{\tilde{v}^{+}}{u^{+}}.$
Then
$\tilde{b}^{-}=\frac{1}{x^{-}u^{-}}\sum_{i\in I_{-}}x_{i}u_{i}\,\frac{v_{i}}{u_{i}},\qquad\tilde{b}^{+}=\frac{1}{x^{+}u^{+}}\sum_{i\in I_{+}}x_{i}u_{i}\,\frac{v_{i}}{u_{i}}.$
Since $x^{-}u^{-}=\sum_{i\in I_{-}}x_{i}u_{i}$ and $\frac{v_{i}}{u_{i}}\in\{b^{-},b^{+}\}$, we conclude that $\tilde{b}^{-}$ is a weighted average of $b^{-}$ and $b^{+}$, and hence $\tilde{b}^{-}\in[b^{-},b^{+}]$. The same argument applies to $\tilde{b}^{+}$ and yields $\tilde{b}^{+}\in[b^{-},b^{+}]$.
Note that in general $\tilde{b}^{\pm}\neq b^{\pm}$ so we cannot simply take $v^{\prime}:=\tilde{v}$ and conclude the proof. An extra step is needed.
Step 2: Monotonicity argument.
For $b_{1},b_{2}>0$ define
$f(b_{1},b_{2}):=F_{2}\!\left(x^{\prime},y^{\prime};u^{\prime},(b_{1}u^{-},\,b_{2}u^{+})\right).$
Then
$f(b_{1},b_{2})=\frac{x^{\prime}\!\cdot u^{\prime}}{y^{\prime}\!\cdot u^{\prime}}\,\frac{Ab_{1}+Bb_{2}}{Cb_{1}+Db_{2}},$
where
$A:=a^{-}\,x^{-}u^{-},\qquad B:=a^{+}\,x^{+}u^{+},\qquad C:=x^{-}u^{-},\qquad D:=x^{+}u^{+}.$
Note that
$AD-BC=x^{-}u^{-}\,x^{+}u^{+}\,(a^{-}-a^{+})\;<\;0,$
since $a^{-}<a^{+}$. Then
$\frac{\partial f}{\partial b_{1}}(b_{1},b_{2})=\frac{x^{\prime}\!\cdot u^{\prime}}{y^{\prime}\!\cdot u^{\prime}}\,\frac{(AD-BC)\,b_{2}}{(Cb_{1}+Db_{2})^{2}}\;<\;0,$
$\frac{\partial f}{\partial b_{2}}(b_{1},b_{2})=\frac{x^{\prime}\!\cdot u^{\prime}}{y^{\prime}\!\cdot u^{\prime}}\,\frac{(BC-AD)\,b_{1}}{(Cb_{1}+Db_{2})^{2}}\;>\;0,$
for all $b_{1},b_{2}>0$. It follows that $f(b_{1},b_{2})$ is decreasing in $b_{1}$ and increasing in $b_{2}$.
Define
$v^{\prime}:=(b^{-}u^{-},\,b^{+}u^{+})\in\mathbb{R}_{>0}^{2},$
so that $s^{\pm}(u^{\prime},v^{\prime})=b^{\pm}$. By construction,
$f(b^{-},b^{+})=F_{2}(x^{\prime},y^{\prime};u^{\prime},v^{\prime})$